### v1

# Sigma: Scalable Spectral Insights for LLM Collapse

Abstract
The rapid adoption of synthetic data for training Large Language Models (LLMs) has introduced the technical challenge of “model collapse”–a degenerative process where recursive training on model-generated content leads to a contraction of distributional variance and representational quality. While the phenomenology of collapse is increasingly evident, rigorous methods to quantify and predict its onset in high-dimensional spaces remain elusive. In this paper, we introduce Sigma (Spectral Inequalities for Gram Matrix Analysis), a unified framework that benchmarks model collapse through the spectral lens of the embedding Gram matrix. By deriving and utilizing deterministic and stochastic bounds on the matrix’s spectrum, Sigma provides a mathematically grounded metric to track the contraction of the representation space. Crucially, our stochastic formulation enables scalable estimation of these bounds, making the framework applicable to large-scale foundation models where full eigendecomposition is intractable. We demonstrate that Sigma effectively captures the transition towards degenerate states, offering both theoretical insights into the mechanics of collapse and a practical, scalable tool for monitoring the health of recursive training pipelines.
1 Introduction
LLM Collapse (shumailov2024; Shumailov2024-dj) refers to the phenomenon where Large Language Models (LLMs) progressively degrade towards a degenerate state. This state arises as a result of recursive training, where models are trained on iterations of data generated by other models, often previous versions of themselves. We term this model-generated output Synthetic Data, in contrast to organic data produced by human users.
In the current landscape of AI research, the demand for reliable and high-quality training data has never been higher. The scale of the latest generation of LLMs is increasing at such a staggering speed that the natural generation of internet content by humans cannot keep pace. Consequently, utilizing synthetic data has become a vital, if not inevitable, step in LLM training. While this is a logical progression, it carries a significant caveat. Without the infusion of new data produced by human users, model collapse occurs as the model iterates on its own probability distribution. Ultimately, the model may converge to a state characterized by low variance and degraded representational quality (shumailov2024; Shumailov2024-dj).
Therefore, it is imperative to conduct a rigorous mathematical examination of LLM collapse. What are the underlying mechanisms at the mathematical level? How can this degradation be tracked and quantified throughout the training process? Furthermore, how can we leverage these metrics to gain deeper insights into LLM dynamics and guide the development of more efficient training architectures? This paper aims to answer these questions.
We introduce Sigma (Spectral Inequalities for Gram Matrix Analysis), a novel framework that quantifies and benchmarks model collapse by utilizing the spectral bounds of the Gram matrix of LLM embedding vectors. The Gram matrix is a classical mathematical concept named after Jørgen Pedersen Gram. It has a wide range of applications in fields such as Riemannian geometry and control theory, and has found critical utility in modern applied mathematics and computer science domains, including the finite element method (10.5555/894170). It is also widely used in classical NLP as an intuitive method to encode variable quantities of natural language embeddings into a fixed-dimensional matrix representation(10.1162/153244302760200687; NIPS2014_b7866697). As we enter the era of LLMs, the Gram matrix remains a vital tool, as it captures the intrinsic geometry and distributional properties of the data we seek to analyze. Sigma revisits this fundamental operator to quantify the modern phenomenon of representation collapse.
Inevitably, the spectrum of the Gram matrix becomes the key to understanding LLM collapse. We develop our theory and framework around this central mathematical object, allowing us to monitor, estimate, and predict the behavior of LLMs regarding model collapse through training iterations. Crucially, our method utilizes both deterministic and stochastic bounds to function effectively even with partially observed data, ensuring our framework remains scalable for large-scale models.
Our contributions can be summarized as follows:
  • A Theoretical Framework for Collapse Detection: We propose Sigma, a spectral analysis framework that rigorously links the decay of the Gram matrix eigenspectrum to the phenomenon of model collapse.
  • Rigorous Spectral Bounds: We derive both deterministic and stochastic inequalities for the Gram matrix, allowing for precise estimation of representation collapse even in high-dimensional settings where exact computation is prohibitively expensive.
  • Scalable Benchmarking Metrics: Leveraging these bounds, we introduce a set of scalable metrics that can track the “health” of an LLM’s distribution during recursive training, providing an early warning system for the onset of degenerate states.
2 Background
2.1 Related Work
The risk of training generative models on their own outputs has been formalized as model collapse, where repeated self-training progressively erodes distributional support, particularly in the tails, and can culminate in highly generic or degenerate generations (shumailov2024; Shumailov2024-dj).
Related phenomena have been studied under different labels and experimental protocols, including Model Autophagy Disorder (MAD), which emphasizes precision–recall tradeoffs and the role of “fresh” real data injections across generations (alemohammad2023selfconsuminggenerativemodelsmad). Recent theory sharpens these concerns by distinguishing regimes where degradation is transient versus strong and asymptotic, even when synthetic data is mixed with a constant fraction of real data (dohmatob2024strongmodelcollapse).
Complementary analyses interpret recursion as an entropy- or variance-contracting process and propose tail-centric quantities such as time-to-forget and covariance shrinkage to explain why rare events vanish first (seddik2024badtrainingsyntheticdata). Other lines formalize stability of self-consuming loops via recursion-stability notions, connecting collapse to unbounded generalization degradation across generations (yoon2025modelcollapseselfconsumingchain).
Finally, practical collapse signals in text include over-concentration of surface features (e.g., $n$-grams), emphasizing the need for diagnostics that are sensitive to diversity loss beyond average quality metrics (zhu2025synthesizetextdatamodel). Most prior work studies recursion within a single-model loop (or a fixed teacher–student chain), whereas real-world synthetic data arises from multi-model, retrieval-mediated ecosystems. Building on this black-box perspective, recent work develops statistically grounded uncertainty and calibration techniques tailored to LLM-as-a-Judge systems (pang2025unsupervisedconformal; pang2025tamingvariability), and studies related dynamics that can shift judge behavior over time (wang2025llmwebdynamicstracing), where LLM Web Dynamics (LWD) explicitly targeting this setting by modeling a network of LLMs coupled through a retrieval-augmented substrate, and analyzes convergence patterns under such web-like feedback.
In contrast to task- or token-level statistics alone, our SIGMA framework focuses on representation geometry: we track collapse via the eigenspectrum of embedding Gram matrices and develop deterministic and stochastic spectral inequalities that enable scalable, one-sided monitoring when only partial Gram information is available.
2.2 Preliminaries
For initial information encoding in LLMs, the $i$-th sentence is mapped to a vector of a fixed dimension $m$. We denote this vector as $X_{i}\in\mathbb{R}^{m}$, which is referred to as the embedding or the embedding vector of the sentence. Let the set of embedding vectors be $E_{0}=\{X_{i}:i\in[n]\}$, where $n]:=\{1,\dots,n\}$. $E_{0}$ represents our starting evaluation corpus. We operate under the regime where the sample size exceeds the embedding dimension, i.e., $n>m$.
Subsequently, we construct an $m\times n$ matrix $M$ by concatenating all embedding vectors column-wise. We define the Gram matrix as:
$G=MM^{\!\top},$ (2.1)
with respect to the set $E_{0}$. Note that $G$ is an $m\times m$ matrix whose dimensions are independent of the sample size $|E_{0}|$, though its entries depend on the samples. Since we assume $n>m$, it is possible for $G$ to be full rank. Given that LLMs are probabilistic models, the set $E$ contains random vectors; consequently, $G$ is treated as a random matrix. Our objective is to establish a benchmark using the spectrum of the Gram matrix to measure the magnitude of model collapse.
It is well known that the Gram matrix $G$ captures the inter-dependency relations between the semantic features of the sentences. We posit that $G$ being full rank is a necessary condition for the LLM to function properly, indicating that the model utilizes the full capacity of its embedding space. Conversely, if $G$ is singular (or effectively rank-deficient), it implies information loss at the semantic level and marks the onset of model collapse.
To provide an intuitive (non-rigorous) example: imagine a collapsed LLM that outputs identical or highly repetitive phrasing for distinct queries. These outputs would map to linearly dependent embedding vectors, failing to span the ambient space $\mathbb{R}^{m}$, thereby resulting in a singular Gram matrix.
Motivated by this, we introduce the log-determinant as our primary benchmark metric for model collapse:
$\log|G|=\log\left(\prod_{j=1}^{m}\lambda_{j}\right)=\sum_{j=1}^{m}\log(\lambda_{j}),$ (2.2)
where $\lambda_{j}$ are the eigenvalues of $G$. When collapse occurs, we expect a sharp decrease in the log-determinant (diverging towards -$\infty$ as $\lambda_{min}\to 0$).
3 The Sigma Framework: Deterministic & Spectral Bounding of the Gram Matrix
To rigorously quantify model collapse, we analyze the spectral properties of the embedding Gram matrix. However, computing the full spectrum for large-scale LLMs is often computationally intractable due to the sheer volume of tokens ($n_{k}$) relative to the embedding dimension ($m$). In this section, we present the Sigma framework, which leverages a sub-sampling strategy to derive both deterministic and stochastic bounds on the Gram determinant. This allows us to estimate the “health” of the representation space using only a fraction of the data.
3.1 Setup and Notations
We track the model’s state at a fixed training iteration $k$. Adding superscripts to the basic notation in (2.2), let $M^{(k)}\in\mathbb{R}^{m\times n_{k}}$ be the embedding matrix, where each column $v^{(k)}_{j}\in\mathbb{R}^{m},1\leq j\leq n_{k}$ is an embedding vector. We define the Gram matrix as:
$G^{(k)}:=M^{(k)}\left(M^{(k)}\right)^{\!\top}=\sum_{j=1}^{n_{k}}v^{(k)}_{j}\left(v^{(k)}_{j}\right)^{\!\top}\in\mathbb{S}^{m}_{+},$ (3.1)
where $\mathbb{S}^{m}_{+}$ denotes the set of $m\times m$ positive semi-definite matrices. We denote the eigenvalues of $G^{(k)}$ by $\lambda_{1}(G^{(k)})\geq\dots\geq\lambda_{m}(G^{(k)})\geq 0$.
The Sub-Sampling Strategy. Since $n_{k}$ is typically massive, we partition the embedding matrix $M^{(k)}$ into two blocks:
$A^{(k)}\in\mathbb{R}^{m\times n_{A}}\text{: The {observed} block;}$
$B^{(k)}\in\mathbb{R}^{m\times n_{B}}\text{: The {unobserved} block.}$
Here, $n_{k}=n_{A}+n_{B}$, and we assume $n_{A}>m$ to ensure rank stability. $A^{(k)}$ consists of the columns we keep during the computation and $B^{(k)}$ contains the part we ignore. Write
$G_{A}^{(k)}=A^{(k)}(A^{(k)})^{\top},\quad G_{B}^{(k)}=B^{(k)}(B^{(k)})^{\top}$ (3.2)
the sub-Gram Matrices formed by $A^{(k)}$ and $B^{(k)}$. A simple computation gives us
$G^{(k)}\ =\ G_{A}^{(k)}+G_{B}^{(k)}.$ (3.3)
Our goal is to bound the spectrum (and determinant) of the full $G^{(k)}$ using only the observable $G_{A}^{(k)}$.
3.2 Main Results
We first provide a rigorous deterministic bound that relies on no distributional assumptions, utilizing Weyl’s inequality and Ky Fan dominance. The detailed deduction as well as the proofs of theorems can be found in appendix A.
Theorem 1 (Deterministic Spectral Bound).
Let $G^{(k)}=G_{A}^{(k)}+G_{B}^{(k)}$ be as defined above. Let $\beta_{k}:=\lambda_{\max}(G_{B}^{(k)})$ be the spectral radius of the unobserved component. Then, the determinant of the full Gram matrix is bounded by:
$\det(G_{A}^{(k)})\ \leq\ \det(G^{(k)})\ \leq\ \prod_{i=1}^{m}\left(\lambda_{i}(G_{A}^{(k)})+\beta_{k}\right).$ (3.4)
While Theorem 1 is rigorous, the dependence on $\beta_{k}$ (the max eigenvalue of the unobserved data) makes it loose in practice. To obtain a scalable estimator for real-world monitoring, we derive a stochastic bound assuming the data is drawn from a consistent underlying distribution. This assumption does not restrict the geometric complexity of the embeddings; it merely posits that the generative process is stationary, ensuring that the spectral properties of the sub-sample generalize to the full corpus.
Theorem 2 (Stochastic Scaling Spectral Bound).
Assume the columns of $M^{(k)}$ are i.i.d. samples drawn from a distribution with covariance matrix $C\in\mathbb{S}^{m}_{+}$. As $n_{A},n_{k}\to\infty$ with fixed $m$, there exists a finite constant $K$ such that with high probability:
$\det(G^{(k)})\ \approx\ \left(\frac{n_{k}}{n_{A}}\right)^{m}\det(G_{A}^{(k)}).$ (3.5)
Specifically, the upper bound holds in expectation:
$\mathbb{E}[\det(G^{(k)})]\ \leq\ K\cdot\left(\frac{n_{k}}{n_{A}}\right)^{m}\det(G_{A}^{(k)}).$ (3.6)
Theorem 2 implies a simple operational scaling law: for a healthy model, the log-determinant increases naturally with sample size as $m\log n_{k}$. This allows us to define a size-invariant baseline: the difference $\log\det G^{(k)}-m\log n_{k}$ should remain constant across checkpoints. By monitoring this corrected value, we isolate genuine geometric collapse from trivial fluctuations in dataset size, ensuring that any observed negative drift reflects a true degradation in the model’s representational capacity.
3.3 Error Estimation
Naturally, we will be interested in how much error does the theorems in the pervious section introduce. In fact, 2 provides extra insights for the tightness of inequalities in 1. Let $C$ be the covariance matirx defined in theorem 2. We know that for large $n_{A},n_{k}$, $\det(G^{(k)})\approx\left(n_{k}/n_{A}\right)^{m}\det(G_{A}^{(k)})\approx n^{m}\det(C)$, we can define the asymptotic Overestimation Ratio $\mathcal{R}$ as the ratio of the Weyl bound to the stochastic expected value:
$\mathcal{R}$ $=\frac{1}{\det G^{(k)}}\prod_{i=1}^{m}\left(\lambda_{i}(G_{A}^{(k)})+\beta_{k}\right)$
$\approx\prod_{j=1}^{m}\frac{n_{A}\lambda_{j}(C)+(n_{k}-n_{A})\lambda_{\max}(C)}{n\lambda_{j}(C)}.$ (3.7)
Dividing through by $n$, and letting $\rho=n_{A}/n$, we have a more concise formula
$\mathcal{R}\approx\prod_{j=1}^{m}\left(\rho+(1-\rho)\frac{\lambda_{\max}(C)}{\lambda_{j}(C)}\right).$ (3.8)
For each $j$, we consider the ratio $\lambda_{\max}(C)/\lambda_{j}(C)\geq 1$. If this ratio is close to $1$ for all $j$, it implies the covariance matrix $C$ is close to being isotropic; consequently, the overestimation ratio will also be close to $1$, yielding an asymptotically tight deterministic bound.
On the other hand, if $C$ is more anisotropic - i.e., the ratio $\lambda_{\max}(C)/\lambda_{j}(C)\gg 1$ for some smaller eigenvalues $\lambda_{j}(C)$ - then the term $(1-\rho)\lambda_{\max}(C)/\lambda_{j}(C)$ will dominate the product in (3.8) and cause the error to grow.
As for theorem 2 itself, we can also give a more refined estimation of the error. Let $X$ be a random vector drawn from the i.i.d. distribution in theorem 2, we set
$\sigma:=\text{Var}\left(X^{T}CX\right).$ (3.9)
The following theorem holds:
Theorem 3 (Stochastic Error Term).
Assume the settings of theorem 1 and, in addition, the distribution of column vectors has a finite fourth moment. Then as $n_{A},n_{k}\to\infty$ with $n_{A}<n_{k}$, we have
$\frac{\log\det G^{(k)}-\log\det G_{A}^{(k)}-m\log(n_{k}/n_{A})}{\sigma\ \sqrt{\frac{1}{n_{A}}-\frac{1}{n_{k}}}}\to\mathcal{N}(0,1)$ (3.10)
in distribution. Consequently, for any $\alpha\in(0,1)$, let $z_{\alpha/2}=\Phi^{-1}(1-\alpha/2)$. Then with probability approaching $1-\alpha$:
$\left|\log\det G_{n}-\log\left(\left(\frac{n_{k}}{n_{A}}\right)^{m}\det G_{A}\right)\right|\quad\quad\quad$
$\leq z_{\alpha/2}\cdot\sigma\sqrt{\frac{1}{n_{A}}-\frac{1}{n_{k}}}.$ (3.11)
Stability in Large Corpora. While theorem 2 establishes the asymptotic scaling law, theorem 3 quantifies the precision of this law in finite samples. Specifically, it provides the exact confidence intervals required to bound the error of our scaling-law estimator in the finite-sample regime. The estimation error depends on the factor $\sqrt{1/n_{A}-1/n_{k}}$. In the practical regime where the total volume of model outputs is massive compared to our observed sample ($n_{k}\gg n_{A}$), this factor simplifies effectively to
$\sqrt{\frac{1}{n_{A}}-\frac{1}{n_{k}}}\;\approx\;\frac{1}{\sqrt{n_{A}}}.$ (3.12)
This leads to a crucial practical insight: the precision of our collapse monitoring depends primarily on the size of our observed block $n_{A}$, not on the total size of the unobserved corpus. Even if the total output $n_{k}$ grows indefinitely, our estimate remains stable.
This motivates the definition of a simple plug-in estimator for the full log-determinant:
$\widehat{L}^{(k)}\;:=\;\log\det G_{A}^{(k)}\;+\;m\log\left(\frac{n_{k}}{n_{A}}\right).$ (3.13)
Using the rigorous result from (3), we derive a practical rule-of-thumb for large datasets ($n_{k}\gg n_{A}$) with $(1-\alpha)$ confidence interval:
$\log\det G^{(k)}\approx\widehat{L}^{(k)}\;\pm\;\frac{z_{\alpha/2}\,\sigma}{\sqrt{n_{A}}}.$ (3.14)
This formula allows practitioners to easily determine the necessary sample size: to achieve a target precision $\varepsilon$, one simply needs to choose an observed block size of $n_{A}\gtrsim(z_{\alpha/2}\,\sigma/\varepsilon)^{2}$.
3.4 Sigma-UB Collapse Monitoring
Section 3.2 establishes the theoretical foundation for detection, but the raw bounds rely on unobserved quantities ($G_{B}^{(k)}$) or unknown constants ($K$). This subsection translates those theorems into two scalar, fully computable diagnostics, computed only from $\left(G_{A}^{(k)},n_{A},n_{k}\right)$.
Regularization Convention ($\delta$). We fix a small constant $\delta>0$ and evaluate log-volumes using $\log\det(\cdot+\delta I_{m})$. The same $\delta$ is reused across all checkpoints to ensure comparability and numerical stability under rank deficiency. This regularization ensures the metric is well-defined and Lipschitz continuous with respect to small perturbations in the embeddings.
Now we turn both theorem 1 and theorem 2 into computationally scalable monitoring metrics via the following two bridges, respectively.
Diagnostic I (Theorem 1): Tail-Energy Budget. Note that we never construct $B^{(k)}$ and thus cannot measure $\beta_{k}$. Instead, we enforce a standard embedding preprocessing (e.g., $\ell_{2}$-normalization or clipping) so that $\|v_{j}^{(k)}\|_{2}^{2}\leq\rho$ for all generated responses. With $n_{2}:=n_{k}-n_{A}$ unseen columns, this yields the computable worst-case budget
$\widehat{\beta}_{k}\;:=\;(n_{k}-n_{A})\,\rho.$ (3.15)
Appendix B shows that substituting $\widehat{\beta}_{k}$ produces a valid one-sided envelope for the full regularized log-determinant.
Diagnostic II (Theorem 2): Constant-Free Size Correction. Recall that Theorem 2 implies the full model entropy is effectively the sub-sample entropy shifted by the geometric scaling factor $m\log(n_{k}/n_{A})$. While the exact value depends on an unknown constant $K$, we can eliminate this constant by monitoring the trend rather than the absolute value.
Combining this with Theorem 1, we operationally track collapse via two complementary diagnostics: (i) a size-corrected statistic that relies on this stochastic scaling, and (ii) a normalized deterministic envelope that uses a computable tail budget.
Two Computable Diagnostics. We monitor collapse using (i) a normalized deterministic envelope and (ii) a covariance-normalized LLN proxy:
$\mathcal{G}_{\mathrm{KF}}^{(k)}(\delta)$ $:=\log\det\!\left(G_{A}^{(k)}+(\widehat{\beta}_{k}+\delta)I_{m}\right)$
$\quad-m\log(\widehat{\beta}_{k}+\delta),$ (3.16)
$\mathcal{U}_{\mathrm{LLN,cov}}^{(k)}(\delta)$ $:=\log\det\!\left(G_{A}^{(k)}+\delta I_{m}\right)-m\log n_{A}.$ (3.17)
We report baseline drifts
$\Delta\mathcal{G}_{\mathrm{KF}}^{(k)}(\delta)$ $:=\mathcal{G}_{\mathrm{KF}}^{(k)}(\delta)-\mathcal{G}_{\mathrm{KF}}^{(0)}(\delta),$
$\Delta\mathcal{U}_{\mathrm{LLN,cov}}^{(k)}(\delta)$ $:=\mathcal{U}_{\mathrm{LLN,cov}}^{(k)}(\delta)-\mathcal{U}_{\mathrm{LLN,cov}}^{(0)}(\delta).$ (3.18)
Sigma-UB Monitoring Algorithm. Algorithm 1 computes both diagnostics at each checkpoint using only the sub-Gram matrix and counts. In practice, $\log\det(\cdot)$ is computed via Cholesky on the shifted PSD matrix (Appendix B).
Algorithm 1 Sigma-UB Collapse Monitoring (upper-bound diagnostics)
[…]
1: Define $\textsc{LogDet}_{\delta}(X):=\log\det(X+\delta I_{m})$.
2:for $k=0,\dots,K$ do
3:  $\widehat{\beta}_{k}\leftarrow(n_{k}-n_{A})\rho$
4:  $S_{k}\leftarrow\textsc{LogDet}_{\delta}(G_{A}^{(k)})$
5:  $\mathcal{G}_{\mathrm{KF}}^{(k)}\leftarrow\textsc{LogDet}_{\delta}(G_{A}^{(k)}+\widehat{\beta}_{k}I_{m})-m\log(\widehat{\beta}_{k}+\delta)$
6:  $\mathcal{U}_{\mathrm{LLN,cov}}^{(k)}\leftarrow S_{k}-m\log n_{A}$
7:  if $k=0$ then
8:   store $(\mathcal{G}_{\mathrm{KF}}^{(0)},\mathcal{U}_{\mathrm{LLN,cov}}^{(0)})$
9:  else
10:   compute drifts $\Delta\mathcal{G}_{\mathrm{KF}}^{(k)}$, $\Delta\mathcal{U}_{\mathrm{LLN,cov}}^{(k)}$ via (3.18)
11:  end if
12:end for
Return: $\Delta\mathcal{G}_{\mathrm{KF}}^{(K)}$, $\Delta\mathcal{U}_{\mathrm{LLN,cov}}^{(K)}$
How to Interpret the Outputs. Both diagnostics decrease when eigenvalues of $G_{A}^{(k)}$ contract, with heightened sensitivity near the $\delta$ floor. Operationally, a sustained negative drift in either $\Delta\mathcal{G}_{\mathrm{KF}}^{(k)}(\delta)$ or $\Delta\mathcal{U}_{\mathrm{LLN,cov}}^{(k)}(\delta)$ signals collapse onset; saturation indicates late-stage rank collapse. The two-track design is intentional: $\Delta\mathcal{G}_{\mathrm{KF}}$ is a deterministic, conservative envelope controlled by a known tail-energy budget, while $\Delta\mathcal{U}_{\mathrm{LLN,cov}}$ is a sensitive size-corrected proxy whose unknown constants cancel under differencing. Disagreement between the tracks is itself informative, often indicating violations of the consistent-context assumption (prompt drift, retrieval drift, filtering changes) rather than numerical artifacts.
4 Experiments
4.1 Experiment Objectives and Controlled Assets
Our experiments measure the operational behavior of training with recursively regenerated synthetic text under tightly controlled, evaluation-stable conditions. We isolate two questions, aligned with two settings (S1–S2):
(1) Data recursion (S1). When synthetic text is regenerated every generation but each training run restarts from the same base checkpoint, do the SIGMA geometry diagnostics drift progressively, indicating representational contraction driven purely by data feedback?
(2) Weight recursion (S2). Under the same regeneration protocol and real:synth mixture rule, does carrying model weights across generations materially amplify or alter collapse dynamics relative to restart-from-base training?
Assets and evaluation stability. We reuse the license-clean real corpora and frozen prompt banks from DecayBench as controlled assets. Each domain
$d\in\{\textsc{CIVICS},\textsc{FINANCE},\textsc{SCIENCE},\textsc{SOFTWARE}\}$
provides: (i) a real corpus $R^{(d)}$ pre-sliced into non-overlapping 64-token contexts, and (ii) a fixed prompt bank $P^{(d)}$ with stable prompt IDs (and bucket labels). To avoid evaluation drift, all checkpoints are evaluated on the same frozen prompt bank $P^{(d)}$ with fixed decoding and a frozen sentence encoder. Following the main-text budget, we report results on a single target domain (FIN) and provide full numeric tables and auxiliary analyses in Appendix C.
4.2 Models and Fine-Tuning Setup
Unless otherwise noted, we start from a single base checkpoint $M^{(0)}$ (e.g., OPT-125M) and fine-tune with a fixed optimizer, training budget, and data formatting across all conditions. All training examples are stored as 64-token text blocks so that token budgets remain comparable. Full hyperparameters and run metadata (optimizer/LR, steps, batch size, weight decay, decoding, seeds) are consolidated in an experiment card (Appendix C.2).
4.3 Synthetic Generation, Mixture Rule, and Two Recursion Settings
Regeneration protocol and constant-budget mixtures. For a fixed domain $d$, we use each real 64-token context $x\in R^{(d)}$ as a prompt and sample a fixed-length continuation from the previous-generation model to form a synthetic pool $S^{(g,d)}$ at generation $g\geq 1$. Training data at generation $g$ is a constant-budget mixture of synthetic and real examples: a fraction $(1-\alpha)$ are drawn from $S^{(g,d)}$ and a fraction $\alpha$ from $R^{(d)}$, with the total number of training blocks held fixed to $|R^{(d)}|$. The real fraction $\alpha\in[0,1]$ is held fixed within each run.
Setting S1: restart-from-base (data recursion only). Each generation regenerates a fresh synthetic pool, but training always restarts from the same base checkpoint:
$M^{(g,d)}\leftarrow\mathrm{Finetune}\!\big(M^{(0)},\,D^{(g,d)}(\alpha)\big).$
This isolates collapse induced by the evolving synthetic data distribution without compounding weight updates across generations.
Setting S2: true recursion (data + weight recursion). This setting carries weights across generations under the same regeneration and mixture rule:
$M^{(g,d)}\leftarrow\mathrm{Finetune}\!\big(M^{(g-1,d)},\,D^{(g,d)}(\alpha)\big).$
Comparing S1 vs. S2 therefore isolates the incremental effect of weight recursion under a matched synthetic regeneration pipeline.
4.4 Collapse Measurements and Baselines
Sigma-UB (two-track geometry diagnostics). At each evaluated checkpoint (generation $g$ in S1/S2), we generate one response per frozen prompt (fixed decoding), embed responses with a frozen sentence encoder, and compute the two SIGMA-UB diagnostics from Section 3.4 (Algorithm 1) on a fixed observed index set (shared across checkpoints within a run). We report baseline drifts relative to the base checkpoint ($g=0$): (i) Track II $\Delta U_{\mathrm{LLN,cov}}(\delta)$ and (ii) Track I $\Delta G_{\mathrm{KF}}(\delta)$.
Interpretation. More negative drift indicates stronger contraction of the observed sub-Gram spectrum. Track II is the sensitive trend probe (observed-spectrum contraction); Track I is a conservative guardrail and may be loose when dominated by the worst-case tail-energy budget. We therefore interpret the joint behavior of both tracks as the primary diagnostic signal.
Surface-form proxies (non-definitional). To contextualize geometry drift, we also compute lightweight surface-form diversity proxies (distinct-2 and hashed $n$-gram HHI) on the same model outputs used for SIGMA-UB. Full proxy reporting is deferred to Appendix C.4.
4.5 Main-Text Reporting: Two Plots and One Table (FIN)
To keep the main text minimal and directly comparable across settings, we report: (i) two plots—one for S1 and one for S2—each showing both SIGMA-UB tracks, and (ii) one compact table summarizing S1 vs. S2 via final drift and an OLS slope over generations. All underlying per-generation values used to render the plots and compute the summary scalars are provided in Appendix C.3.
Figure 1: FIN, Setting S1 (restart-from-base): drift vs. generation under both SIGMA-UB tracks. Top: Track II $\Delta U_{\mathrm{LLN,cov}}(\delta)$. Bottom: Track I $\Delta G_{\mathrm{KF}}(\delta)$. Values are baseline drifts relative to the base checkpoint ($g=0$). S1 isolates data recursion only: synthetic text is regenerated each generation, but training restarts from the same base checkpoint. Appendix C.3 reports the exact per-generation table used to render this figure.
Figure 2: FIN, Setting S2 (true recursion): drift vs. generation under both SIGMA-UB tracks. Top: Track II $\Delta U_{\mathrm{LLN,cov}}(\delta)$. Bottom: Track I $\Delta G_{\mathrm{KF}}(\delta)$. Values are baseline drifts relative to the base checkpoint ($g=0$). S2 compounds data recursion + weight recursion: each generation regenerates synthetic text and continues training from the previous generation weights. Appendix C.3 reports the exact per-generation table used to render this figure.
Setting $g_{\max}$ $\Delta U_{\mathrm{LLN,cov}}$ (final) slopeLLN,cov $\Delta G_{\mathrm{KF}}$ (final) slopeKF Track agreement?
S1 50 -150.986 -0.941 -142.051 -0.918 yes (both negative)
S2 50 -1536.562 -42.621 -2.426 2.350 no (Track II collapses; Track I does not)
Table 1: FIN summary statistics (S1 vs. S2). We summarize each drift curve using: (i) final drift at the last available generation $g_{\max}$, and (ii) an ordinary least-squares slope of drift versus generation (more negative implies faster degradation over the observed window; slopes are fit on the full drift series including the g=0g{=}0 baseline point). Track agreement reports whether both tracks trend negative (qualitative agreement) or whether Track II indicates strong contraction without a matching Track I response. Appendix C.3 provides the per-generation tables used to compute these scalars.
Result (S1 vs. S2). On FIN, restart-from-base recursion (S1) shows a mild but consistently negative drift under both tracks by $g_{\max}=50$ (final drifts $\approx-151$ and $\approx-142$), consistent with gradual contraction driven by the evolving synthetic pool. Under true recursion (S2), Track II exhibits a much larger negative drift (final $\approx-1537$; slope $\approx-42.6$), indicating substantially accelerated contraction when weights are carried across generations. In contrast, Track I remains near-zero and can even trend slightly positive over the same window, reflecting that the KF envelope is governed by a worst-case tail-energy budget rather than the observed spectrum itself; this Track II/Track I separation is an intended diagnostic feature of Sigma-UB.
Acknowledgements
The first author would like to thank his friend Xiaoyu Li for his unwavering support throughout the difficult times.
Appendix A Mathematical Deductions and Proof of Theorems
We Adopt the notations from section 3.1, for the Gram matrix $G^{(k)}$, if $\mathrm{rank}(M^{(k)})<m$ (e.g., if $n_{k}<m$ or the columns of $M^{(k)}$ are linearly dependent), then $G^{(k)}$ is singular and $\det(G^{(k)})=0$. In practice, we usually have $n_{k}\gg m$, and the Cauchy–Binet formula for Gram matrices yields
$\det(G^{(k)})\ =\ \sum_{J\subset[n_{k}],\,|J|=m}\ \det(M^{(k)}_{J})^{2},$ (A.1)
where $M^{(k)}_{J}$ is the $m\times m$ submatrix formed by columns indexed by $J$. Recall from Section 2.2 that computing the full Gram matrix can be computationally prohibitively expensive. Instead, we compute the log-determinant of the sub-Gram matrix, formed by selecting a subset of the columns in the embedding matrix.
Proof of Theorem 1.
To prove the lower bound, notice that $A^{(k)}$ is a sub-matrix of $M^{(k)}$. By the Cauchy Binet formula (A.1), the sum defining $\det(G_{A}^{(k)})$ is a subset of the sum defining $\det(G^{(k)})$. This directly implies
$\det(G_{A}^{(k)})\leq$ $\det(G^{(k)}).$ (A.2)
For the second half of the inequality, we use Weyl’s inequality and Ky Fan dominance and for each $i$:
$\lambda_{i}(G^{(k)})\ \leq\ \lambda_{i}(G_{A}^{(k)})+\lambda_{1}(G_{B}^{(k)}).$ (A.3)
Hence, with $\beta_{k}:=\lambda_{\max}(G_{B}^{(k)})$, we have:
$G^{(k)}\preceq$ $\ G_{A}^{(k)}+\beta_{k}I_{m},\quad\det(G^{(k)})\leq\prod_{i=1}^{m}\left(\lambda_{i}(G_{A}^{(k)})+\beta_{k}\right).$ (A.4)
Combining the lower bound in (A.2) and the upper bound in (A.4) yields the stated result. ∎
The Cauchy-Binet formula, as well as the bound in Eq. (A.4), opens a window for us to bound the determinant of the original Gram matrix. Alas, the upper bound is not sharp because there is no way for us to extract information from the columns that we do not choose.
However, in practice, this is usually not the case because the dataset we are working on typically comes from a consistent context. If we use the analogy of viewing the Gram matrix as a book, then the chapters we read (i.e., $G_{A}^{(k)}$) are usually a good indication of the chapters we do not read (i.e., $G_{B}^{(k)}$), even though mathematically the unseen chapters could contain content that is completely irrelevant.
To reflect such realistic circumstances, we assume that the embedding column vectors are drawn from an implicit distribution (probability measure) $\mu$. Denote $v$ as an arbitrary column (where $v$ can be seen as a random variable) and let $C=\mathbb{E}[vv^{\top}]$ be the covariance matrix. With this additional assumption, We proceed with the asymptotic argument.
Proof of Theorem 2.
By the Strong Law of Large Numbers,
$\frac{1}{n_{k}}G^{(k)}\to C\quad\text{a.s.}$ (A.5)
For the original Gram matrix, since $n_{k}$ is usually very large, we have
$\det(G^{(k)})\approx\det(n_{k}C)=n_{k}^{m}\det(C).$ (A.6)
On the other hand, assuming $n_{A}$ is sufficiently large to approximate the population covariance, then
$\det(G_{A}^{(k)})\approx O(\det(n_{A}C))=n_{A}^{m}\det(C).$ (A.7)
Therefore, we have a stochastic upper bound of the original Gram matrix using the sub-Gram matrix: there exists a finite constant $K<\infty$ such that for each iteration $k$,
$\det(G^{(k)})\leq K\cdot\left(\frac{n_{k}}{n_{A}}\right)^{m}\det(G_{A}^{(k)}).$ (A.8)
This bound serves as an asymptotic scaling law rather than a point-wise estimator of $\det(G^{(k)})$. It gives us a theoretical foundation for monitoring model collapse with partial data, assuming the bare minimum: that the data we are processing comes from a consistent context.
Proof of Theorem 3.
Since we assume that the distribution of column vectors have a finite fourth moment, we know
$Z:=X^{T}C^{-1}X=\sum_{i,j=1}^{m}X_{i}X_{j}$ (A.9)
has a finite variance. i.e., $\sigma$ is finite. In fact, $Z$ is the Mahalanobis distance. It plays an important role in this proof. For lighter notation, we write $n\equiv n_{k}$ and $G_{t}\equiv G_{t}^{(k)}$ the gram matrix formed by selecting $t$ columns from $M^{(k)}$. This way $G^{(k)}\equiv G_{n}$ and $G^{(k)}_{A}\equiv G_{n_{A}}$. Notice that
$\det(G_{t+1})=\det(G_{t})(1+X_{t+1}^{T}G_{t}^{-1}X_{t+1}).$ (A.10)
Denote $L_{n}=\log\det G_{n}$ and we get
$L_{n}-L_{n_{A}}=\sum_{t=n_{A}}^{n-1}\log\left(1+X_{t+1}^{T}G_{t}^{-1}X_{t+1}\right).$ (A.11)
Write the normalized deviation matrix as $\Delta_{t}:=tG_{t}^{-1}-C^{-1}$. We need the following lemma:
Lemma 4 (Convergence Rate of Deviation Matrix).
Under the assumption of finite fourth moments, the operator norm of the deviation matrix satisfies $\|\Delta_{t}\|_{op}=O_{p}(t^{-1/2})$.
Proof of Lemma 4.
Let $\hat{C}_{t}=\frac{1}{t}G_{t}=\frac{1}{t}\sum_{i=1}^{t}X_{i}X_{i}^{T}$. Since the random vectors $X_{i}$ are i.i.d. with finite fourth moments, the elements of the matrix product $X_{i}X_{i}^{T}$ have finite variances. Applying the Central Limit Theorem to the vectorized matrix, we find that $\sqrt{t}(\hat{C}_{t}-C)$ converges in distribution to a zero-mean Gaussian random matrix. Consequently, the Euclidean norm (and thus the operator norm) of the difference is bounded in probability:
$\lVert\hat{C}_{t}-C\rVert_{op}=O_{p}(t^{-1/2})$ (A.12)
Then, by the resolvent identity for invertible matrices $A$ and $B$, which states $A^{-1}-B^{-1}=-A^{-1}(A-B)B^{-1}$. Setting $A=\hat{C}_{t}$ and $B=C$, we obtain
$\Delta_{t}=\hat{C}_{t}^{-1}-C^{-1}=-\hat{C}_{t}^{-1}(\hat{C}_{t}-C)C^{-1}.$ (A.13)
Taking the operator norm of both sides:
$\lVert\Delta_{t}\rVert_{op}\leq\lVert\hat{C}_{t}^{-1}\rVert_{op}\lVert\hat{C}_{t}-C\rVert_{op}\lVert C^{-1}\rVert_{op}.$ (A.14)
By the Weak Law of Large Numbers, $\hat{C}_{t}\to C$ in probability. Since matrix inversion is a continuous function on the manifold of positive definite matrices, $\|\hat{C}_{t}^{-1}\|_{op}\to\|C^{-1}\|_{op}$ in probability as well, by the continuous mapping theorem. Therefore, $\|\hat{C}_{t}^{-1}\|_{op}=O_{p}(1)$. Substituting the rates gives us
$\|\Delta_{t}\|_{op}=O_{p}(1)\cdot O_{p}(t^{-1/2})\cdot O(1)=O_{p}(t^{-1/2}).$ (A.15)
With the above lemma being stated and proved, we rewrite the quadratic form exactly as
$X_{t+1}^{T}G_{t}^{-1}X_{t+1}=\frac{1}{t}X_{t+1}^{T}(C^{-1}+\Delta_{t})X_{t+1}=\frac{1}{t}Z_{t+1}+\frac{1}{t}R_{t+1},$ (A.16)
where $Z_{t+1}=X_{t+1}^{T}C^{-1}X_{t+1}$ is the squared Mahalanobis distance, and $R_{t+1}=X_{t+1}^{T}\Delta_{t}X_{t+1}$ is the remainder.
Then, using the Taylor expansion $\log(1+u)=u-\frac{u^{2}}{2}+O(u^{3})$, valid for small $u$, yields
$\log\left(1+\frac{Z_{t+1}+R_{t+1}}{t}\right)=\frac{Z_{t+1}}{t}+\frac{R_{t+1}}{t}-\frac{1}{2}\left(\frac{Z_{t+1}}{t}\right)^{2}+O_{p}\left(\frac{1}{t^{3}}\right).$ (A.17)
We define $\xi_{t+1}=Z_{t+1}-m$. Note that $\mathbb{E}(Z_{t+1})=\text{Tr}(C^{-1}C)=m$. Thus $\mathbb{E}(\xi_{t+1})=0$ and recall from (3.9) that $\sigma^{2}:=\text{Var}(Z_{t+1})$. Substituting $Z_{t+1}=m+\xi_{t+1}$:
$\log\left(1+\frac{Z_{t+1}+R_{t+1}}{t}\right)=\frac{m+\xi_{t+1}}{t}+\frac{R_{t+1}}{t}-\frac{m^{2}+2m\xi_{t+1}+\xi_{t+1}^{2}}{2t^{2}}+O_{p}\left(\frac{1}{t^{3}}\right).$ (A.18)
Rearranging the terms by the order of $t$:
$\log\left(1+\frac{Z_{t+1}+R_{t+1}}{t}\right)=\frac{m}{t}+\frac{\xi_{t+1}}{t}+\underbrace{\frac{R_{t+1}}{t}-\frac{m^{2}}{2t^{2}}+O_{p}\left(\frac{1}{t^{2.5}}\right)}_{\text{Higher Order Terms}}.$ (A.19)
Summing from $n_{A}$ to n1n-1 gives us the decomposition
$L_{n}-L_{n_{A}}=\sum_{t=n_{A}}^{n-1}\frac{m}{t}+\sum_{t=n_{A}}^{n-1}\frac{\xi_{t+1}}{t}+\mathcal{E}_{n,n_{A}},$ (A.20)
where $\mathcal{E}_{n,n_{A}}$ is the cumulative error formed by summing over all the higher order terms in (A.19).
To prove the theorem, we remain to examine each component in (A.20). For $\mathcal{E}_{n,n_{A}}$, by lemma 4, $R_{t+1}=O_{p}(t^{-1/2})$ and thus $R_{t+1}/t=O_{p}(t^{-3/2})$. The summation of $R_{t+1}$ becomes the tail of an absolutely convergent series. Furthermore, standard bias corrections imply this term decays faster than the leading stochastic fluctuation located in the second term, rendering it asymptotically negligible. The entire $\mathcal{E}_{n,n_{A}}$ will vanish.
For the first summation, using the Euler-Maclaurin expansion $\sum_{t=a}^{b}\frac{1}{t}=\ln(b)-\ln(a)+O(a^{-1})$ and we have
$\sum_{t=n_{A}}^{n-1}\frac{m}{t}=m\ln(n/n_{A})+O(n_{A}^{-1}).$ (A.21)
Finally, for the second summation, the summands $\xi_{t+1}/t$ are independent, centered random variables with variance
$\text{Var}\left(\frac{\xi_{t+1}}{t}\right)=\frac{\sigma^{2}}{t^{2}}.$ (A.22)
The variance of the summation satisfies as $n,n_{A}\to\infty$,
$\text{Var}\left(\sum_{t=n_{A}}^{n-1}\left(\frac{\xi_{t+1}}{t}\right)\right)=\sum_{t=n_{A}}^{n-1}\frac{\sigma^{2}}{t^{2}}\to\sum_{t=n_{A}}^{n-1}\frac{\sigma^{2}}{t(t+1)}=\sigma^{2}\left(\frac{1}{n_{A}}-\frac{1}{n}\right).$ (A.23)
Combining (A.20), (A.21), (A.23) and the fact $\mathcal{E}_{n,n_{A}}$ vanishes, we have, by the Central Limit Theorem,
$\frac{\log\det G_{n}-\log\det G_{n_{A}}-m\log(n/n_{A})}{\sigma\ \sqrt{\frac{1}{n_{A}}-\frac{1}{n}}}\to\mathcal{N}(0,1)$ (A.24)
in distribution, as desired. ∎
Appendix B Mathematical Details for Sigma-UB(Section 3.4)
This appendix contains the derivations that connect the theoretical bounds in Section 3.2 to the computable diagnostics in Algorithm 1.
B.1 Regularized $\log$-determinant and PSD Monotonicity
Fix $\delta>0$. For $X,Y\in\mathbb{S}^{m}_{+}$, if $X\preceq Y$ then $X+\delta I_{m}\preceq Y+\delta I_{m}$ and
$\log\det(X+\delta I_{m})\leq\log\det(Y+\delta I_{m}),$ (B.1)
since all eigenvalues are strictly positive after the $\delta I_{m}$ shift.
B.2 Diagnostic I: from $\beta_{k}$ to a Computable Budget $\widehat{\beta}_{k}$
Since $G_{B}^{(k)}\succeq 0$,
$\beta_{k}:=\lambda_{\max}(G_{B}^{(k)})\leq\mathrm{tr}(G_{B}^{(k)}).$ (B.2)
Using $G_{B}^{(k)}=\sum_{j\in B}v_{j}^{(k)}(v_{j}^{(k)})^{\top}$ gives
$\mathrm{tr}(G_{B}^{(k)})=\sum_{j\in B}\|v_{j}^{(k)}\|_{2}^{2}.$ (B.3)
Assume preprocessing enforces $\|v_{j}^{(k)}\|_{2}^{2}\leq\rho$ for all columns. With $n_{2}:=n_{k}-n_{A}$ unseen columns,
$\beta_{k}\leq\widehat{\beta}_{k}:=n_{2}\rho.$ (B.4)
The above energy cap and tail budget gives us the PSD Majorant (Deterministic Envelope). Indeed, from $G^{(k)}=G_{A}^{(k)}+G_{B}^{(k)}$ and Ky-Fan/Weyl dominance,
$G^{(k)}\preceq G_{A}^{(k)}+\beta_{k}I_{m}\preceq G_{A}^{(k)}+\widehat{\beta}_{k}I_{m}.$ (B.5)
Shifting by $\delta I_{m}$ and applying PSD monotonicity yields the valid envelope gives us
$\log\det(G^{(k)}+\delta I_{m})\leq\log\det\!\left(G_{A}^{(k)}+(\widehat{\beta}_{k}+\delta)I_{m}\right)=:\mathcal{U}_{\mathrm{KF}}^{(k)}(\delta).$ (B.6)
Now let $\mathcal{U}^{(k)}_{\mathrm{KF},\star}(\delta):=\log\det(G_{A}^{(k)}+(\beta_{k}+\delta)I_{m})$. Because $\widehat{\beta}_{k}\geq\beta_{k}$, we obtain
$0\leq\mathcal{U}_{\mathrm{KF}}^{(k)}(\delta)-\mathcal{U}^{(k)}_{\mathrm{KF},\star}(\delta)=\sum_{i=1}^{m}\log\!\left(\frac{\lambda_{i}(G_{A}^{(k)})+\widehat{\beta}_{k}+\delta}{\lambda_{i}(G_{A}^{(k)})+\beta_{k}+\delta}\right)\leq m\log\!\left(\frac{\widehat{\beta}_{k}+\delta}{\beta_{k}+\delta}\right).$ (B.7)
Rearrange the above inequality and we arrive at the conceptual normalization inequality, which tells us the the gap is controlled by the ideal (incomputable) envelope.
$\mathcal{U}_{\mathrm{KF}}^{(k)}(\delta)-m\log(\widehat{\beta}_{k}+\delta)\leq\mathcal{U}^{(k)}_{\mathrm{KF},\star}(\delta)-m\log(\beta_{k}+\delta).$ (B.8)
Motivated by the deduction above, we can define the normalized reporting form as
$\mathcal{G}_{\mathrm{KF}}^{(k)}(\delta)$ $:=\mathcal{U}_{\mathrm{KF}}^{(k)}(\delta)-m\log(\widehat{\beta}_{k}+\delta),$ (B.9)
which matches (3.16).
B.3 Diagnostic II: LLN Scaling, Cancellation of $K$, and Covariance Normalization
From the LLN scaling route, we can set
$\mathcal{U}_{\mathrm{LLN}}^{(k)}(\delta):=\log\det(G_{A}^{(k)}+\delta I_{m})+m\log\!\left(\frac{n_{k}}{n_{A}}\right).$ (B.10)
If $\log\det(G^{(k)})\leq\log K+\mathcal{U}_{\mathrm{LLN}}^{(k)}(0)$ for an unknown finite $K$, then baseline differencing removes $\log K$:
$\Delta\mathcal{U}_{\mathrm{LLN}}^{(k)}(\delta):=\mathcal{U}_{\mathrm{LLN}}^{(k)}(\delta)-\mathcal{U}_{\mathrm{LLN}}^{(0)}(\delta).$ (B.11)
This means that we effectively remove the unknown constant $K$ from the diagnostic. Finally, we can define the covariance-normalized form
$\mathcal{U}_{\mathrm{LLN,cov}}^{(k)}(\delta):=\mathcal{U}_{\mathrm{LLN}}^{(k)}(\delta)-m\log n_{k}=\log\det(G_{A}^{(k)}+\delta I_{m})-m\log n_{A},$ (B.12)
which is (3.17) and is the statistic computed in Algorithm 1.
B.4 Numerical Computation of $\log$-determinant
For $S\succ 0$, compute $\log\det(S)$ via Cholesky $S=LL^{\top}$:
$\log\det(S)=2\sum_{i=1}^{m}\log L_{ii}.$ (B.13)
Apply this to $G_{A}^{(k)}+\delta I_{m}$ and $G_{A}^{(k)}+(\widehat{\beta}_{k}+\delta)I_{m}$.
Appendix C Supplementary Experiments, Reproducibility, and Extended Diagnostics (FIN)
C.1 Section Roadmap and Artifact Integrity
This section of the appendix serves three purposes.
(i) Reproducibility.
We provide a complete specification of the experimental pipeline (training, synthetic regeneration, evaluation, and Sigma-UB computation) in a set of ten self-contained tables (Appendix C.2). The intent is that an external group can rerun the entire loop without relying on implicit conventions.
(ii) Plot–table consistency.
[…]
(iii) Additional diagnostics and interpretation.
Because Sigma-UB is a geometric diagnostic (not a surface-form metric), we also report lightweight surface-form proxies—distinct-2 and hashed $n$-gram HHI—computed on the same model outputs used for Sigma-UB. We include a compact overlay plot that normalizes all signals to $[0,1]$ to visually assess co-movement (Appendix C.4). We then add a bucket-localization slice (optional) to show where contraction is most pronounced under the frozen prompt taxonomy (Appendix C.5). Throughout, we emphasize that surface-form proxies are contextual and not collapse definitions.
Note on filenames. For continuity with earlier drafts, some exported artifact filenames retain the prefix tech_ (e.g., tech_s1_two_tracks.pdf). In this appendix and in the main text, these files correspond to the FIN run described in Section 4.
C.2 Complete reproducibility specification for S1 and S2
We summarize every operational choice that materially affects outcomes: model hyperparameters; synthetic regeneration and mixture; decoding; seeds; Sigma-UB computation; evaluation metrics; hardware/software; file layout; and the formal difference between S1 and S2. Unless explicitly stated, S1 and S2 share the same configuration; the only intended difference is whether weights are reset to $M^{(0)}$ (S1) or carried forward $M^{(g-1)}$ (S2).
Component S1 S2 Notes
Base model facebook/opt-125m facebook/opt-125m 125M parameters
Fine-tuning epochs / gen 5 5 Fixed per-generation budget
Batch size 16 16 Effective batch size held constant
Learning rate $2\times 10^{-5}$ $2\times 10^{-5}$ Constant across generations
Optimizer AdamW AdamW Weight decay $0.01$
Precision FP16 enabled FP16 enabled GPU training
Table 2: Table 1: Model and training hyperparameters. These hyperparameters are held fixed across S1 and S2 so that observed differences are attributable to data recursion vs. data+weight recursion, rather than to optimization budget changes.
Item Value Applies to Notes
Real fraction $\alpha$ $0.0$ (pure synthetic) S1, S2 Stress-tests recursion (no real-data anchoring)
Context unit 64 tokens S1, S2 Training blocks; constant token budget
Real corpus size $\sim$1,000–1,200 contexts / domain S1, S2 Domain-dependent (DecayBench)
Prompt bank size $\sim$163–200 prompts / domain Eval Frozen prompt IDs and bucket labels
Generations reported $g=0,\dots,50$ (51 checkpoints) Main text FIN run shown in Figures 12
Synthetic pool size $|S^{(g)}|=|R|$ S1, S2 One synthetic continuation per real context per generation
Table 3: Table 2: Data generation and mixing. We use constant-budget mixtures and, for the reported runs, set $\alpha=0$ to isolate synthetic recursion dynamics without real-data refresh.
Decoding parameter Value Used for Notes
Sampling type Nucleus (top-$p$) Gen + Eval Stochastic sampling (not greedy)
Temperature 1.0 Gen + Eval Fixed across generations
Top-pp 0.95 Gen + Eval Fixed across generations
Max tokens (training) 64 Corpus construction 64-token blocks for training
Max tokens (evaluation) 256 Prompt-bank responses Used to obtain sufficiently long responses for embedding/geometry
Table 4: Table 3: Text generation (decoding) parameters. Decoding is frozen across checkpoints to prevent evaluation drift and to keep the generation distribution comparable over time.
Aspect Specification
Base seed 42
Domain offsets Domain-specific offsets are derived via a deterministic hash (to avoid accidental seed reuse across domains).
Per-context determinism Synthetic generations are seeded deterministically per real-context ID and generation index.
Per-sample determinism When multiple sampling steps exist (e.g., batching), the seed schedule ensures reproducible ordering and reproducible stochastic draws.
Implementation note Seed utility functions are included in the codebase; reruns with the same config reproduce identical synthetic pools and identical evaluation outputs.
Table 5: Table 4: Seeds and determinism. The recursion loop is extremely sensitive to small sampling changes; therefore we enforce determinism at the level of domain, generation, and context ID.
Sigma-UB component Value Used for Notes
Sentence encoder
sentence-transformers/
all-MiniLM-L6-v2
Embedding Frozen encoder, $m=384$
Regularization $\delta$ $10^{-3}$ Track I/II Matches Section 3.4 conventions
Energy cap $\rho$ $1.0$ Track I Worst-case tail-energy budget
Baseline $g=0$ Drift computation $\Delta(\cdot)$ reported relative to base checkpoint
Observed block fixed $\mathcal{I}_{A}$ within run Track I/II Prevents measurement drift across checkpoints
Table 6: Table 5: Sigma-UB computation. Both tracks are computed from the same frozen embeddings and the same observed index set, so drift is attributable to model evolution rather than to evaluation sampling.
Metric family Reported quantity Notes
Sigma-UB Track II $U_{\mathrm{LLN,cov}}$ and $\Delta U_{\mathrm{LLN,cov}}$ Sensitive trend probe (observed-spectrum contraction)
Sigma-UB Track I $G_{\mathrm{KF}}$ and $\Delta G_{\mathrm{KF}}$ Conservative envelope; can be loose or non-monotone
Surface proxies distinct-2, hashed $n$-gram HHI Computed on the same generated outputs as Sigma-UB
Excluded metrics No perplexity We intentionally do not report PPL in this revision
Table 7: Table 6: Evaluation metrics. The appendix focuses on geometry diagnostics and lightweight surface-form proxies; perplexity-based baselines are excluded by design.
Component Specification
Training hardware NVIDIA GPU (server environment)
Metrics hardware CPU (Mac M1 used for aggregation/plotting)
Core libraries PyTorch 2.0+, Transformers 4.30+, sentence-transformers 2.2+
Precision Mixed precision (FP16) enabled during training
Table 8: Table 7: Computational environment. Hardware does not change the definition of metrics but can affect throughput; determinism is enforced via fixed seeds and fixed decoding.
Artifact Path template / convention
S1 metrics CSV <METRICS_ROOT>/final/<DOMAIN>/metrics_paper*.csv
S2 metrics CSV <METRICS_ROOT>/true_recursion/<DOMAIN>/ metrics_paper_recursion*.csv
Bucket metrics CSV (optional) <METRICS_ROOT>/final/<DOMAIN>/ metrics_buckets_paper*.csv
Main-text figures figures/<prefix>_s1_two_tracks.pdf, figures/<prefix>_s2_two_tracks.pdf
Point tables tables/<prefix>_s1_two_tracks_points.tex (and S2 analog)
Summary table tables/tab_<prefix>_s1s2_summary.tex
Surface proxy plot figures/<prefix>_surface_proxies.pdf
Table 9: Table 8: Data paths and file structure. This table documents the exact expected locations of input metrics and the produced figures/tables used by the paper.
Setting Defining update rule
S1 (restart-from-base) $M^{(g,d)}\leftarrow\mathrm{Finetune}\!\big(M^{(0)},\,D^{(g,d)}(\alpha)\big)$. Weights do not compound across generations; only the synthetic data distribution evolves.
S2 (true recursion) $M^{(g,d)}\leftarrow\mathrm{Finetune}\!\big(M^{(g-1,d)},\,D^{(g,d)}(\alpha)\big)$. Both synthetic data and weights compound across generations.
Table 10: Table 9: Key difference between S1 and S2. This is the core experimental intervention: the regeneration protocol and mixture rule are held fixed while weight carryover is toggled.
Prompt bucket Description (frozen taxonomy)
Creative Open-ended creative prompts emphasizing novelty and stylistic variation
Divergent Prompts encouraging multiple alternatives, brainstorming, or broad exploration
Analogy / ELI5 Prompts asking for analogies, simplifications, or explanatory reframings
What-if Counterfactual or perturbation prompts (“what if we change X?”)
Neutral Factual/neutral prompts with minimal stylistic pressure
Table 11: Table 10: Prompt buckets for localization. Bucket IDs and assignments are fixed by DecayBench; slicing by bucket does not change the evaluation distribution over time, it only reveals where contraction is concentrated within a frozen taxonomy.
C.3 Exact per-generation values for the main-text drift plots
Figures 12 are generated from exported per-generation tables. We include them here to (a) make the plots auditable and (b) guarantee that all reported summary scalars (final drift and OLS slope) can be recomputed from a single canonical source. In particular, the S1 vs. S2 summary in Table 1 is computed directly from these per-generation series (including the baseline point $g=0$, where drift is 0 by definition).
Generation $g$ $\Delta U_{\mathrm{LLN,cov}}(\delta)$ (Track II) $\Delta G_{\mathrm{KF}}(\delta)$ (Track I)
0 0.000000 0.000000
1 -50.937706 -42.771322
2 -87.220752 -78.285600
3 -89.193112 -80.257959
4 -108.940054 -100.004901
5 -138.689852 -130.138891
6 -135.146696 -126.211544
7 -144.505214 -135.954253
8 -158.731260 -149.796107
9 -151.887859 -142.952706
10 -149.844314 -140.909162
11 -171.990854 -163.055701
12 -181.172379 -172.237226
13 -143.036471 -134.101319
14 -146.954698 -138.019546
15 -150.025009 -141.089857
16 -169.047237 -160.112084
17 -134.754288 -125.819136
18 -116.984750 -108.049597
19 -136.991334 -128.056181
20 -135.678684 -126.743532
21 -137.759104 -128.823951
22 -150.287784 -141.352631
23 -154.278070 -145.342917
24 -132.792859 -123.857707
25 -108.036399 -99.101246
Table 12: Per-generation values for Figure 1 (S1; restart-from-base). Track II and Track I are reported as drifts relative to the base checkpoint ($g=0$). In S1, both tracks remain consistently negative at the end of the run ($g=50$), indicating gradual contraction driven by the evolving synthetic pool even without weight carryover.
Table 13: Per-generation values for Figure 2 (S2; true recursion). Track II reaches large-magnitude negative drift (e.g., beyond 1000-1000 by roughly $g\approx 20$ in this run), indicating rapid observed-spectrum contraction under weight carryover. Track I remains near zero and can be non-monotone (including occasional positive excursions), consistent with its role as a conservative envelope controlled by a tail-energy budget rather than by the observed spectrum itself.
Interpreting Track-I/Track-II disagreement.
A recurring pattern in S2 is that Track II indicates strong contraction while Track I does not “collapse” in the same way. This is not a contradiction: Track II measures observed-spectrum log-volume contraction (trend-sensitive), while Track I is an upper-envelope controlled by a worst-case tail-energy budget. When the tail-energy bound is loose relative to the observed sub-Gram spectrum, Track I can be non-alarming even when Track II contracts sharply. In practice, we treat the separation itself as informative: it flags “early collapse” regimes where the observed spectrum has contracted but the worst-case envelope remains permissive.
C.4 Surface-form proxy diagnostics and alignment with Sigma-UB
Sigma-UB is explicitly designed to track representation geometry rather than surface repetition. Nevertheless, it is useful to ask whether geometry contraction coincides with familiar surface-form failure modes. We therefore compute two lightweight proxies on the same generated text used for Sigma-UB embeddings: (i) distinct-2 (unique/total bigrams; higher is more diverse), and (ii) hashed $n$-gram HHI (higher is more concentrated/repetitive). These proxies are not collapse definitions and can be noisy, but they provide an operational sanity check.
Figure 3: Normalized overlay of Sigma-UB drifts and surface-form proxies (FIN). Each curve is min–max normalized to $[0,1]$ within its panel for visual comparability. We overlay four signals across generations: Track II drift $\Delta U_{\mathrm{LLN,cov}}(\delta)$, Track I drift $\Delta G_{\mathrm{KF}}(\delta)$, distinct-2, and hashed $n$-gram HHI. Left: S2 (true recursion). Right: S1 (restart-from-base). The key qualitative trend is that the large early Track-II contraction in S2 precedes (and is accompanied by) stronger surface-form concentration (rising HHI) than in S1.
What the overlay reveals.
Figure 3 makes two practical points. First, in S2 the Track-II drift becomes strongly negative early, whereas hashed HHI rises more gradually and can spike later. This indicates that Sigma-UB Track II can act as an earlier indicator of contraction than surface repetition alone. Second, Track I remains near the bottom of the normalized scale in S2 for much of the run, reinforcing that Track I is not intended as a sensitive trend detector; it is a conservative envelope that can remain loose even under clear contraction signals elsewhere.
Endpoint comparison and co-movement.
To quantify the qualitative overlay, Table 14 reports proxy endpoints ($g=0$ vs. $g=50$) and simple correlation diagnostics between Track II drift and the proxies over generations. The key operational takeaway is that true recursion (S2) yields substantially larger surface-form concentration (HHI increases by nearly three orders of magnitude relative to its $g=0$ baseline) than restart-from-base (S1), and this increase co-varies strongly with Track II under a rank-based measure.
[…]
S1 0.731 0.334 0.000102 0.013385 0.926 -0.286 -0.318
S2 0.715 0.254 0.000100 0.097262 -0.165 -0.468 -0.685
Table 14: Surface-proxy alignment with Track II (FIN). Distinct-2 decreases and hashed $n$-gram HHI increases in both settings, with substantially stronger concentration under true recursion (S2). Correlations are computed over generations using the per-generation series (including $g=0$). In S1, $\Delta U$ and distinct-2 co-move strongly (both decrease), yielding a high positive Pearson correlation. In S2, distinct-2 is noisier (and can transiently increase despite strong contraction), while HHI retains a strong monotone association with $\Delta U$ (Spearman $\approx-0.69$), consistent with the overlay in Figure 3.
Full per-generation proxy table (optional).
For completeness and auditability, the full per-generation proxy values can be included as a table. This is often useful when debugging changes to tokenization, hashing, or evaluation decoding.
C.5 Within-domain localization by frozen prompt buckets (optional)
Because the evaluation prompt bank is labeled by prompt type (creative, divergent, analogy/ELI5, what-if, neutral), we can localize contraction to specific regimes while keeping the evaluation distribution fixed over time. Concretely, we re-run the evaluation pipeline but restrict to prompts in a single bucket, generating one response per prompt under the same decoding, embedding with the same frozen encoder, and computing the same Sigma-UB drifts. This is a slice of a fixed evaluation set, not a moving target, so it does not introduce evaluation drift.
Figure 4: Bucket localization under S1 (FIN; optional). Sigma-UB drifts computed after restricting evaluation to each frozen prompt bucket. Bucket curves can be noisier than the aggregate because each bucket contains fewer prompts; we therefore interpret bucket localization as a qualitative “where does collapse concentrate?” tool rather than as a primary metric.
Table 15: Track-II drift by bucket under S1 (FIN; optional). Each entry is $\Delta U_{\mathrm{LLN,cov}}(\delta)$ computed using only prompts from the given bucket (baseline $g=0$). In this run, Creative tends to exhibit the largest-magnitude negative drift by late generations, while other buckets are comparatively less contracted. Because bucket sizes differ, cross-bucket comparisons should be interpreted qualitatively unless normalized for prompt count.
Interpretation.
Bucket localization helps answer a practical question: is contraction uniformly distributed across prompt types, or concentrated where the model is pushed into higher-entropy regimes? In the reported S1 run, the Creative bucket shows consistently larger-magnitude Track-II contraction than the aggregate, suggesting that open-ended prompts amplify the representational narrowing induced by synthetic recursion. That said, bucket series are noisier (and can be non-monotone) because each bucket contains fewer prompts and fewer embedded blocks, so we use them as a diagnostic slice rather than as a headline number.